OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x)=(x-8)/(x+7) and g(x)=(-7x-8)/(x-1)

OpenStudy (anonymous):

no one wants to do this algebra, because it is a real pain

OpenStudy (anonymous):

That's not the right attitude to get a problem done now is it? @satellite73

OpenStudy (anonymous):

have fun

OpenStudy (anonymous):

\[f(x)=\frac{x-8}{x+7}\] \[g(x)=\frac{-7x-8}{x-1}\]

OpenStudy (anonymous):

\[f\circ g(x)=f\left(\frac{-7x-8}{x-1}\right)\] \[=\frac{\frac{-7x-8}{x-1}-8}{\frac{-7x-8}{x-1}+7}\]

OpenStudy (anonymous):

clear the compound fraction by multiplying top and bottom by \(x-1\) use parentheses and do the algebra carefully

OpenStudy (anonymous):

when you do it there will be an orgy of cancellation the likes of which is usually unseen and you will be left with only \(x\)

OpenStudy (anonymous):

Thank you for your help @satellite73 I appreciate it, have a good night.

OpenStudy (anonymous):

you too good luck with this one